OpenStudy (anonymous):

6.03] The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? 330 510 520 660

OpenStudy (anonymous):

...x - 1-step y - 2-step x + y = 840 3x - 2y = 870 --Use substitution-- 3(840 - y) - 2y = 870 --Solve for y-- y = 330 --plug y in the first equation-- x + 330 = 840 --solve for x-- x = 510